40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 25 15:43:29 2016, ericq, Update, SUS, OSEM issues SRM_OSEMS.pngPRM_LR.png
    Reply  Thu Aug 25 20:07:35 2016, gautam, Update, SUS, OSEM issues - maybe resolved? 
    Reply  Tue Sep 6 18:21:13 2016, rana, Update, SUS, OSEM issues PRMswing.png
       Reply  Tue Sep 6 20:52:42 2016, gautam, Update, SUS, OSEM issues P9060254.JPG
          Reply  Wed Sep 7 17:23:26 2016, gautam, Update, SUS, PRM LR fixed for now IMG_3208.JPGIMG_3211.JPG
             Reply  Tue Oct 11 00:43:50 2016, ericq, Update, SUS, PRM LR problematic again PRM_LR_Problem.pngsince_fix.pnglast10.png
                Reply  Tue Oct 11 10:15:04 2016, Steve, Update, SUS, PRM LR problematic again PRMalfoiled.png
                   Reply  Tue Oct 11 13:30:49 2016, gautam, Update, SUS, PRM LR problematic again PRM_LR_decay.png
                      Reply  Wed Oct 12 13:34:28 2016, gautam, Update, SUS, PRM LR problematic again 08.png
Message ID: 12476     Entry time: Wed Sep 7 17:23:26 2016     In reply to: 12475     Reply to this: 12546
Author: gautam 
Type: Update 
Category: SUS 
Subject: PRM LR fixed for now 

[johannes, gautam]

  • We took the heavy door off in the morning with Steve's help
  • The problem was quickly identified as the Al foil on the back of the PRM OSEMs (placed to mitigate scattered light making it into the OSEM that was making locking difficult) shorting out the pins on the rear of the OSEM
  • We decided against using a black glass beam stop behind the PRM - rather, we decided to go for Al foil hats that were
    1. More "domed" - so the back plane of the OSEM isn't in direct contact with the Al foil, though the hats themselves are secure and shouldn't simply fly off during pump down etc
    2. Have a piece of kapton (courtesy Koji from the OMC lab) in the dome so that even if the foil hats move around slightly, there should be no danger of accidentally shorting out any pins
  • Without removing the PRM OSEMs, we were only able to image UR and UL unambiguously showing that they have no filters. Not a single of the 5 spare 'short' OSEMs have filters. We have to open the ITMY chamber to reposition the OSEMs in the near future, which is when we will inspect SRM for filters.
  • Attachment #1 shows a picture of these foil hats - the ones actually put on are shaped slightly differently, but the idea is the same
  • Attachment #2 shows the PRM with its new OSEM hats (we also used a piece of clean copper wire to tie the OSEM cables to the tower on the bottom left of the cage as viewed from the BS-PRM chamber door)
  • After closing up the BS-PRM chamber, I locked the IMC to see if the input pointing had gone way off because of our work on the table and the reputation of the tip-tilts hysteresis - I can see weak flashes in the Y arm but not enough to lock - so I will tweak the alignment a little
  • Once I can recover the Y arm alignment, we can move on to peeling first contact and putting the X arm optics in.

Edit 7.30pm: I have managed to recover Y-arm in air locking, and the transmission is up at ~0.6 again which is what we were seeing prior to touching anything on the BS-PRM table, so it looks like the tip-tilt has not gone badly astray... I have also restored the Satellite boxes so that both PRM and SRM have their designated boxes

Attachment 1: IMG_3208.JPG  5.559 MB  | Hide | Hide all
IMG_3208.JPG
Attachment 2: IMG_3211.JPG  5.668 MB  | Hide | Hide all
IMG_3211.JPG
ELOG V3.1.3-